OpenStudy (anonymous):

Consider the functions f(x) = 4 sin(πx) g(x) = 4 sin(πx) − 1 Which statement best describes the relationship between the graphs of f(x) and g(x)? 1. g is a vertical shift of f upward by 1 units. 2. g is a horizontal shift of f to the left by 1 units. 3. g is a vertical shift of f downward by 1 units. 4. None of these. 5. g is a horizontal shift of f to the right by 1 units

hero (hero):

Vertical shift downward by 1

hero (hero):

These are really easy

OpenStudy (anonymous):

i dont understand! explain please

hero (hero):

Hmmm, I'm not good at explaining things

hero (hero):

Basically, it's the graph of the function at every point except you shift it down by -1 at every point

hero (hero):

So ultimately the whole f(x) shifts down by -1 all at the same time when you graph it

hero (hero):

Thus creating g(x)

OpenStudy (anonymous):

so if it is +1 then it shift upward?

hero (hero):

yes, if +1 is outside the parentheses

OpenStudy (anonymous):

how do you know if its vertical or horizontal?

OpenStudy (anonymous):

think about a list of numbers, inputs and outputs. in the first you put in a bunch of numbers, get out a bunch of numbers. in the second you put in the same numbers, get out the same numbers and then subtract 1. so outputs are all one less than they used to be. that shifts everything down one unit

OpenStudy (anonymous):

so if i have \[f(x)\] and \[f(x)+3\] i know \[f(x)+3\] is identical to f(x) except shifted up three units

OpenStudy (anonymous):

ooh very simple

OpenStudy (anonymous):

yes it is simple. it is a bit different when you compare \[f(x)\text { and } f(x+2)\] however

OpenStudy (anonymous):

in this case you add two before you evaluate the function. this has the effect of shifting the graph HORIZONTALLY two units LEFT

OpenStudy (anonymous):

compare \[f(x)=x^2\] with \[f(x)=(x+2)^2\] and you will see why it is shifted left

OpenStudy (anonymous):

ok this one is shifting horizontally but not upward nor downward